OpenStudy (pujan20):

I need help with HW: Which of the following is equal to the square root of the cube root of 6? 6^1/6 6^1/3 6^2/3 6^3/2 I got 6^2/3 but im unsure. I don't want JUST the answer though

jhonyy9 (jhonyy9):

how you think sqrt(a) = a^(1/2) yes ? cuberoot(a)=a^(1/3) sqrt(cuberoot(a)) = ?

OpenStudy (pujan20):

Yea i think i got it wrong. I finsihed the assignment anyways :p

jhonyy9 (jhonyy9):

but now you know what will be the right answer ?